OpenStudy (bbb911):

Add or subtract. Then simplify by collecting like radical terms, if possible. Assume that all expressions under radicals represent nonnegative numbers.

OpenStudy (bbb911):

\[7\sqrt[4]{6} + 6\sqrt[3]{6}\]

OpenStudy (bbb911):

heres a photo

OpenStudy (bbb911):

did u see

rishavraj (rishavraj):

yup...solving the same..are u havng options

OpenStudy (bbb911):

no :0

rishavraj (rishavraj):

\[its~ 7*6^{1/4} + 6^{4/3}\]

OpenStudy (bbb911):

u multipy them?

rishavraj (rishavraj):

multiply what?? its given 7*6^(1/4)

OpenStudy (bbb911):

cos none of my answers have that

rishavraj (rishavraj):

\[write~\sqrt[3]{6} = 6^{1/3}~and~then~multiply~\it~from~6~\]

OpenStudy (bbb911):

oh i seee thank you !

rishavraj (rishavraj):

u r most welcome